OpenStudy (bananas1234):

The general equation of a plane is Ax + By + Cz = D, where A, B, C, and D are real numbers and A is nonnegative. Find the equation of the plane containing the points (2, 0, 0), (0, 4, 0), and (0, 0, 7). Show each step of your process. Then graph the plane.

OpenStudy (bananas1234):

@zepdrix

OpenStudy (bananas1234):

@Nnesha

OpenStudy (bananas1234):

@jim_thompson5910

OpenStudy (bananas1234):

@mathmale

jimthompson5910 (jim_thompson5910):

What topics has your teacher gone over so far? Have you learned about vectors and cross products?

OpenStudy (bananas1234):

I don't remember learning about that, but this is actually an assignment i didn't do from a few months ago. I'm doing it now to get a few extra points. I forgot most of the information, so that's why I need help. So it is possible but it doesn't ring a bell.

jimthompson5910 (jim_thompson5910):

ok one moment

jimthompson5910 (jim_thompson5910):

if we had something like (2, 0, 0) then this means (x,y,z) = (2, 0, 0) and x = 2 y = 0 z = 0 do you agree?

OpenStudy (bananas1234):

yes that makes sence

jimthompson5910 (jim_thompson5910):

Let's plug in (x,y,z) = (2, 0, 0) Ax + By + Cz = D A*2 + B*0 + C*0 = D 2A = D D = 2A

jimthompson5910 (jim_thompson5910):

What do you get when you follow those same basic steps, but plug in (x,y,z) = (0,4,0) ?

OpenStudy (bananas1234):

Ax + By + Cz = D A*0 + B*2 + C*0 = D D= 0A??

jimthompson5910 (jim_thompson5910):

y = 4, not y = 2

OpenStudy (bananas1234):

OK

jimthompson5910 (jim_thompson5910):

so hopefully you can see how D = 4B ?

OpenStudy (bananas1234):

yes i plugged in the wrong number

jimthompson5910 (jim_thompson5910):

ok what happens when you plug in (x,y,z) = (0,0,7) ?

OpenStudy (bananas1234):

Ax + By + Cz = D A*0 + B*0 + C*7 = D D = C7?

jimthompson5910 (jim_thompson5910):

or D = 7C

jimthompson5910 (jim_thompson5910):

So to recap, so far we know... D = 2A D = 4B D = 7C agreed?

OpenStudy (bananas1234):

yep :)

jimthompson5910 (jim_thompson5910):

D is equal to 2A, 4B, 7C there are infinitely many values for D. The smallest whole number D can be is the LCM of 2,4,7 what is the lowest common multiple of 2,4, and 7?

OpenStudy (bananas1234):

it would be 2^1 2^2 and 7^1 so 28?

jimthompson5910 (jim_thompson5910):

yep, so D = 28 which allows us to find the values of A,B,C

OpenStudy (bananas1234):

ok cool

OpenStudy (bananas1234):

so now would have 28 is equal to 2A, 4B, 7C. right?

jimthompson5910 (jim_thompson5910):

if 28 = 2A, then A = ???

OpenStudy (bananas1234):

umm would you divide however many 2's go into 28?

jimthompson5910 (jim_thompson5910):

yes divide both sides by 2

OpenStudy (bananas1234):

A=14

jimthompson5910 (jim_thompson5910):

yep, how about B?

OpenStudy (bananas1234):

7

jimthompson5910 (jim_thompson5910):

and finally C?

OpenStudy (bananas1234):

4

jimthompson5910 (jim_thompson5910):

summary A = 14 B = 7 C = 4 D = 28 making the equation to be Ax + By + Cz = D 14x + 7y + 4z = 28

OpenStudy (bananas1234):

ok awesome! So now how do i graph it?

jimthompson5910 (jim_thompson5910):

are you allowed to use technology like a graphing program?

OpenStudy (bananas1234):

Yeah i think that's fine

jimthompson5910 (jim_thompson5910):

first solve `14x + 7y + 4z = 28` for z to get 14x + 7y + 4z = 28 4z = 28-14x-7y z = (28-14x-7y)/4 now onto the graph there are many options to go with. Google is one good choice. If you were to google `z = (28-14x-7y)/4` it will show you what the graph looks like http://prntscr.com/b3tzd0 the graph is interactive which means you can rotate things around at any angle

jimthompson5910 (jim_thompson5910):

The only issue is that the x,y, and z axis aren't labeled too well.

jimthompson5910 (jim_thompson5910):

you can also just graph by using the x, y, and z intercepts plot the 3 points, then draw a plane through them all

OpenStudy (bananas1234):

Ok i will give it a try

OpenStudy (bananas1234):

This is what i was given to graph on

jimthompson5910 (jim_thompson5910):

ok so you'll plot the three points like so. See attached

OpenStudy (bananas1234):

Ok thank you so much for all your help!!

jimthompson5910 (jim_thompson5910):

imagine a flat plane going through all the points the plane is shown as the purple triangle. Imagine it extended out infinitely in all directions along the flat surface (which is tilted up)

jimthompson5910 (jim_thompson5910):

if I drew the plane with each side extended out infinitely, then it would fill up the entire space making it look non-tilted, but it is tilted the interactive plotter shows what I mean

OpenStudy (bananas1234):

ok that makes sense